The administration fee at the fair is $1.50 for children and $4.00 for adults.On a certain day 2200 people enter the fair and $5050 is collected.Write a system of equation and then solve it.
Answer by Cromlix(4381)

Hi there,
Make x = children
y = adults.
x + y = 2200........(1)
1.50x + 4y = 5050...(2)
Multiply Eq(1) by 1.50
1.50x + 1.50y = 3300.....(1)
1.50x + 4y = 5050........(2)
Subtract Eq (1) from (2)
2.5y = 1750
y = 700
Substitute y = 700 into
Equation (1)
x + 700 = 2200
x = 2200 - 700
x = 1500
No. of children = 1500
No. of adults = 700
Hope this helps :-)